# Robustness-Inspired Defense Against Backdoor Attacks on Graph Neural Networks

### v1

@@ Appendix B Additional Empirical Results on Prediction Variance Caused by Edge Dropping
In this section, we show additional results on prediction variance caused by dropping trigger edges and clean edges on Cora and PubMed datasets. We generate 40 triggers for Cora and 160 triggers for PubMed. The attack method used is DPGBA zhang2024rethinking . The model architecture is a 2-layer GCN kipf2016semi . The results are shown in Fig. 4. From the figure, it is evident that dropping adversarial edge connecting backdoor trigger will typically result in a much larger prediction variance than dropping clean edges.
@@ Appendix C Time Complexity Analysis
Following blakely2021time , to show the computation of a GCN, first we define $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}$, the adjacency matrix with self-loops. $\hat{\mathbf{D}}$ is the diagonal degree matrix of $\hat{\mathbf{A}}$. $\mathbf{A}^{\prime}$ is the normalized adjacency matrix given as:
@@
Thus, we analyze the complexity of the forward step by decomposing Equation (8) into three high-level operations:
@@ Appendix D Detailed Proofs
Assumptions on Graphs. Following dai2022label ; ma2021homophily , we consider a graph $\mathcal{G}$, where each node $v_{i}$ has features $\mathbf{x}_{i}\in\mathbb{R}^{m}$, a label $y_{i}$ and $\operatorname{deg}(i)$ denotes the number of its neighbors. We assume that: (1) The feature of node $v_{i}$ is sampled from the feature distribution $F_{y_{i}}$ that depends on the label $y_{i}$ of the node, i.e., $\mathbf{x}_{i}\sim\mathcal{F}_{y_{i}}$. This means that the feature of the node is influenced by its label, with $\mu(\mathcal{F}_{y_{i}})$ denoting the mean of the distribution. (2) Feature dimensions of $\mathbf{x}_{i}$ are independent to each other; (3) The features in $\mathbf{X}$ are bounded by a positive scalar $B$, i.e., $\max_{i,j}|\mathbf{X}[i,j]|\leq B$;
@@ D.1 Proof of Theorem 1
Consider a graph $\mathcal{G}=\left\{\mathcal{V},\mathcal{E},\mathbf{X}\right\}$ following Assumptions (1)-(3). For clean node $v_{i}\in\mathcal{V}$ and its neighbors $\mathcal{N}(i)$, the expectation of the pre-activation output of a single operation defined in Eq. (3) is given by $\mathbb{E}[\mathbf{h}_{i}]$. Then the expectation of the pre-activation output of a single operation defined in Eq. (3) after random edge dropping is given by $\mathbb{E}[\mathbf{h}_{i}^{k}]=\mathbb{E}[\mathbf{h}_{i}]$.
Proof. Given the graph convolution operation defined in Eq. (3), we obtain node representation $\mathbb{E}[\mathbf{h}_{i}]=\mathbb{E}\left[\sum_{j\in\mathcal{N}(i)}\frac{1}{\sqrt{deg(i)}\sqrt{deg(j)}}\mathbf{W}\mathbf{x}_{j}\right]$ For any node $v_{i}\in\mathcal{V}$ and its neighbors $\mathcal{N}(i)$, after random edge dropping with drop ratio $\beta$, in expectation, each neighbor in $\mathcal{N}(i)$ contributes with a weight of $(1-\beta)$, and the expected $\operatorname{deg}(i)$ is also scaled by $(1-\beta)$. Therefore, the expectation of the pre-activation output of a single operation defined in Eq. (3) after random edge dropping is given by
@@ D.2 Proof of Theorem 2
$\small\mathbb{P}\left(\left\|\mathbf{h}_{i}^{k}-\mathbb{E}\left[\mathbf{h}_{i}\right]\right\|_{2}\geq t\right)\leq 2\cdot M\cdot\exp\left(-\frac{\operatorname{deg}(i)_{k}t^{2}}{2\rho^{2}(\mathbf{W})B^{2}M}\right),$ (10)
@@
Proof. Following dai2022label ; ma2021homophily , we consider a $d$-regular graph $\mathcal{G}$, i.e. each node has $d$ neighbors, then the expectation of $\mathbf{h}_{i}$ can be derived as $\mathbb{E}[\mathbf{h}_{i}]=\mathbb{E}\left[\sum_{j\in\mathcal{N}(i)}\frac{1}{deg(i)}\mathbf{W}\mathbf{x}_{j}\right]$. Let $\mathbf{x}_{j}^{k}[m],n=1,\dots,M$ denote the $m$-th element of $\mathbf{x}_{j}^{k}$. Then, for any dimension $m$, $\{x_{j}^{k}[m],j\in\mathcal{N}(i)\}$ is a set of independent random variables bounded by $[-B,B]$. Hence, directly applying Hoeffding’s inequality, for any $t_{1}\geq 0$, we have the following bound:
$\small\mathbb{P}\left(\left|\frac{1}{\mathcal{N}(i)^{k}}\sum_{j\in\mathcal{N}(i)^{k}}\left(\mathbf{x}_{j}^{k}[m]-\mathbb{E}\left[\mathbf{x}_{j}^{k}[m]\right]\right)\right|\geq t_{1}\right)\leq 2\exp\left(-\frac{(\operatorname{deg}(i))_{k}t_{1}^{2}}{2B^{2}}\right)$ (11)
@@ D.3 Proof of Theorem 3
To empirically verify that our method of random edge dropping can efficiently distinguish poisoned nodes from clean nodes, we conduct experiments on the Cora Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 , PubMed, and OGB-arxiv NEURIPS2020_fb60d411 datasets, using 40 triggers for Cora, 160 triggers for PubMed and 565 triggers for OGB-arxiv, respectively. The attack method used is DPGBA zhang2024rethinking . The model architecture is a 2-layer GCN kipf2016semi . We set the drop ratio $\beta=0.5$ and performed $K=10$ iterations of random edge dropping. The results are shown in Fig. 5. In Figs. 5(a) and (b), we show the results for Cora and OGB-arxiv, respectively. In Fig. 5(c), we show the result for OGB-arxiv with two edges linking a backdoor trigger and a target node. Our observations are as follows: (i) Our method consistently results in higher prediction variance for poisoned nodes, thus enabling the distinction between poisoned nodes and clean nodes. (ii) Even when two edges link a backdoor trigger to a poisoned node, our method still results in high prediction variance for most of the poisoned nodes. This demonstrates the superior performance of our method compared to the intuitive approach of dropping each edge individually. Additional visualizations of prediction variance under different attacks on various datasets are in Fig. 6, Fig. 7 and Fig 8.
@@
Appendix F Training Algorithm
We summarize the training method of RIGBD for training a backdoor robust GNN node classifier in Algorithm 1. Specifically, we start by randomly initializing the parameters $\theta_{b}$ for an L-layer GNN node classifier $f_{b}$, which uses the graph convolution operation defined in Eq. (3) (line 1). In each iteration of the training loop (lines 2-4), we update $\theta_{b}$ by training $f_{b}$ on the backdoored graph $\mathcal{G}_{T}$ using supervised learning until the model converges. After convergence, we enter a second loop (lines 5-8) where we conduct random edge dropping on $\mathcal{G}_{T}$ and perform inference with the backdoor model $f_{b}$ (line 6). For each node, we calculate the prediction variance $s(i)$ using Eq. (2) (line 7). Nodes with prediction variance $s(i)>\tau$ are selected as candidate poisoned nodes $\mathcal{V}_{s}$, where $\tau$ is calculated by Eq. (5). Next, we randomly initialize $\theta$ for another L-layer GNN node classifier $f$ (line 10). In the final training loop (lines 11-13), we update $\theta$ by training $f$ using the robust training strategy defined in Eq. (6) until convergence. The algorithm concludes by returning the backdoor robust GNN node classifier $f$ (line 14).
@@ Appendix F Training Algorithm
1:  Randomly initialize $\theta_{b}$ for a $L$-layer GNN node classifier $f_{b}$ with graph convolution operation defined in Eq. (3);
@@
7:     Calculate the prediction variance $s(i)$ for each node by Eq. (2);
@@
9:  Select candidate poisoned nodes $\mathcal{V}_{s}$ for nodes with prediction variance $s(i)>\tau$, where $\tau$ is calculated by Eq. (5).
@@
12:     Update $\theta$ by training $f$ using the robust training strategy defined in Eq. (6);
@@
Appendix G Additional Details of Experiment Settings
G.1 Dataset Statistics
Cora and PubMed are citation networks where nodes denote papers, and edges depict citation relationships. In Cora, each node is described using a binary word vector, indicating the presence or absence of a corresponding word from a predefined dictionary. In contrast, PubMed employs a TF/IDF weighted word vector for each node. For both datasets, nodes are categorized based on their respective research areas. OGB-arxiv is a citation network encompassing all Computer Science arXiv papers cataloged in the Microsoft Academic Graph. Each node is characterized by a 128-dimensional feature vector, which is derived by averaging the skip-gram word embeddings present in its title and abstract. Additionally, the nodes are categorized based on their respective research areas. The statistics details of these datasets are summarized in Table 5.
@@ G.2 Attack Methods
    GTA xi2021graph : GTA utilizes a trigger generator that crafts subgraphs as triggers tailored to individual samples. The optimization of the trigger generator focuses exclusively on the backdoor attack loss, disregarding any constraints related to trigger detectability.
@@
    UGBA Dai_2023 : UGBA selects representative and diverse nodes as poisoned nodes to fully utilize the attack budget. An adaptive trigger generator is optimized with a constraint loss so that the generated triggers are ensured to be similar to the target nodes.
@@
    DPGBA Dai_2023 : DPGBA introduce adversarial learning strategy to generate in-distribution triggers. A novel loss is proposed to help adaptive trigger generator to generate efficient in-distribution triggers.
G.3 Compared Methods
@@
    Prune Dai_2023 : Prune is designed to prune edges that linking nodes with low similarity to remove those triggers that are totally different from clean nodes.
@@
    OD zhang2024rethinking : OD introduces graph auto-encoder and filter out nodes with high reconstruction loss to remove those triggers that are outliers compared to clean ones.
@@
    ABL li2021anti : ABL exploits two key weaknesses of backdoor attacks: 1) models learn backdoored data much faster than clean data, with stronger attacks causing faster convergence; 2) the backdoor task is tied to a specific class. Thus, they propose a local gradient ascent (LGA) technique to trap the loss value of each example around a certain threshold. They also propose a global gradient ascent (GGA) loss function to unlearn the backdoor with a small subset of backdoor examples while continuing to learn from the remaining clean examples.
@@
    RS wang2021certified : Randomized smoothing on graphs was first proposed to counter adversarial structural perturbations. We adopt this method and set the drop ratio $\beta>0.5$ to balance the defense performance and clean accuracy. To further ensure clean accuracy, we introduce an adversarial training strategy during the training phase.
@@
    GNNGuard zhang2020gnnguard : GNNGuard is a robust defense method for GNNs that protects against adversarial attacks by leveraging node similarity to filter out adversarial edges. It employs a multi-stages defense strategy that dynamically adjusts edge weights during training, enhancing the model’s resilience to structural perturbations.
@@
    RobustGCN zhu2019robust : RobustGCN enhances the robustness of GCN against adversarial attacks by using Gaussian distributions as node representations, which absorb the effects of adversarial changes. It also introduces a variance-based attention mechanism that assigns different weights to node neighborhoods based on their variances, reducing the propagation of adversarial effects through the graph.
G.4 Implementation Details

### v2

@@ Appendix B Additional Empirical Results on Prediction Variance Caused by Edge Dropping
In this section, we show additional results on prediction variance caused by dropping trigger edges and clean edges on Cora and PubMed datasets. We generate 40 triggers for Cora and 160 triggers for PubMed. The attack method used is DPGBA (Zhang et al., 2024). The model architecture is a 2-layer GCN (Kipf & Welling, 2016). The results are shown in Fig. 4. From the figure, it is evident that dropping adversarial edge connecting backdoor trigger will typically result in a much larger prediction variance than dropping clean edges.
@@ Appendix C Time Complexity Analysis
Following (Blakely et al., 2021), to show the computation of a GCN, first we define $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}$, the adjacency matrix with self-loops. $\hat{\mathbf{D}}$ is the diagonal degree matrix of $\hat{\mathbf{A}}$. $\mathbf{A}^{\prime}$ is the normalized adjacency matrix given as:
@@
Thus, we analyze the complexity of the forward step by decomposing Equation 8 into three high-level operations:
@@ Appendix D Detailed Proofs
Assumptions on Graphs. Following (Dai et al., 2022b; Ma et al., 2021a), we consider a graph $\mathcal{G}$, where each node $v_{i}$ has features $\mathbf{x}_{i}\in\mathbb{R}^{m}$, a label $y_{i}$ and $\operatorname{deg}(i)$ denotes the number of its neighbors. We assume that: (1) The feature of node $v_{i}$ is sampled from the feature distribution $F_{y_{i}}$ that depends on the label $y_{i}$ of the node, i.e., $\mathbf{x}_{i}\sim\mathcal{F}_{y_{i}}$. This means that the feature of the node is influenced by its label, with $\mu(\mathcal{F}_{y_{i}})$ denoting the mean of the distribution. (2) Feature dimensions of $\mathbf{x}_{i}$ are independent to each other; (3) The features in $\mathbf{X}$ are bounded by a positive scalar $B$, i.e., $\max_{i,j}|\mathbf{X}[i,j]|\leq B$;
@@ D.1 Proof of Theorem 1
Consider a graph $\mathcal{G}=\left\{\mathcal{V},\mathcal{E},\mathbf{X}\right\}$ following Assumptions (1)-(3). For clean node $v_{i}\in\mathcal{V}$ and its neighbors $\mathcal{N}(i)$, the expectation of the pre-activation output of a single operation defined in Eq. 3 is given by $\mathbb{E}[\mathbf{h}_{i}]$. Then the expectation of the pre-activation output of a single operation defined in Eq. 3 after random edge dropping is given by $\mathbb{E}[\mathbf{h}_{i}^{k}]=\mathbb{E}[\mathbf{h}_{i}]$.
Proof. Given the graph convolution operation defined in Eq. 3, we obtain node representation $\mathbb{E}[\mathbf{h}_{i}]=\mathbb{E}\left[\sum_{j\in\mathcal{N}(i)}\frac{1}{\sqrt{deg(i)}\sqrt{deg(j)}}\mathbf{W}\mathbf{x}_{j}\right]$ For any node $v_{i}\in\mathcal{V}$ and its neighbors $\mathcal{N}(i)$, after random edge dropping with drop ratio $\beta$, in expectation, each neighbor in $\mathcal{N}(i)$ contributes with a weight of $(1-\beta)$, and the expected $\operatorname{deg}(i)$ is also scaled by $(1-\beta)$. Therefore, the expectation of the pre-activation output of a single operation defined in Eq. 3 after random edge dropping is given by
@@ D.2 Proof of Theorem 2
$\mathbb{P}\left(\left\|\mathbf{h}_{i}^{k}-\mathbb{E}\left[\mathbf{h}_{i}\right]\right\|_{2}\geq t\right)\leq 2\cdot M\cdot\exp\left(-\frac{\operatorname{deg}(i)_{k}t^{2}}{2\rho^{2}(\mathbf{W})B^{2}M}\right),$ (10)
@@
Proof. Following (Dai et al., 2022b; Ma et al., 2021a), we consider a $d$-regular graph $\mathcal{G}$, i.e. each node has $d$ neighbors, then the expectation of $\mathbf{h}_{i}$ can be derived as $\mathbb{E}[\mathbf{h}_{i}]=\mathbb{E}\left[\sum_{j\in\mathcal{N}(i)}\frac{1}{deg(i)}\mathbf{W}\mathbf{x}_{j}\right]$. Let $\mathbf{x}_{j}^{k}[m],n=1,\dots,M$ denote the $m$-th element of $\mathbf{x}_{j}^{k}$. Then, for any dimension $m$, $\{x_{j}^{k}[m],j\in\mathcal{N}(i)\}$ is a set of independent random variables bounded by $[-B,B]$. Hence, directly applying Hoeffding’s inequality, for any $t_{1}\geq 0$, we have the following bound:
$\mathbb{P}\left(\left|\frac{1}{\mathcal{N}(i)^{k}}\sum_{j\in\mathcal{N}(i)^{k}}\left(\mathbf{x}_{j}^{k}[m]-\mathbb{E}\left[\mathbf{x}_{j}^{k}[m]\right]\right)\right|\geq t_{1}\right)\leq 2\exp\left(-\frac{(\operatorname{deg}(i))_{k}t_{1}^{2}}{2B^{2}}\right)$ (11)
@@ D.3 Proof of Theorem 3
To empirically verify that our method of random edge dropping can efficiently distinguish poisoned nodes from clean nodes, we conduct experiments on the Cora (Sen et al., 2008), PubMed, and OGB-arxiv (Hu et al., 2020) datasets, using 40 triggers for Cora, 160 triggers for PubMed and 565 triggers for OGB-arxiv, respectively. The attack method used is DPGBA (Zhang et al., 2024). The model architecture is a 2-layer GCN (Kipf & Welling, 2016). We set the drop ratio $\beta=0.5$ and performed $K=10$ iterations of random edge dropping. The results are shown in Fig. 5. In Figs. 5(a) and (b), we show the results for Cora and OGB-arxiv, respectively. In Fig. 5(c), we show the result for OGB-arxiv with two edges linking a backdoor trigger and a target node. Our observations are as follows: (i) Our method consistently results in higher prediction variance for poisoned nodes, thus enabling the distinction between poisoned nodes and clean nodes. (ii) Even when two edges link a backdoor trigger to a poisoned node, our method still results in high prediction variance for most of the poisoned nodes. This demonstrates the superior performance of our method compared to the intuitive approach of dropping each edge individually. Additional visualizations of prediction variance under different attacks on various datasets are in Fig. 6, Fig. 7 and Fig 8.
@@
Appendix F Empirical Results of Prediction Variance for Nodes with Low Degrees and Poisoned Nodes
In this section, to address concerns about high prediction variance in low-degree or heterophily nodes after random edge dropping, we compare the prediction variance of nodes from both homophily and heterophily graphs against poisoned nodes. We utilize the UGBA attack method to generate backdoor triggers for each dataset, setting $K$ to 10 and the drop ratio to 0.5. For each dataset, we select nodes with degrees 1, 2, and 3. Our experiments are conducted on homophily graphs (Cora, Citeseer, Pubmed, Physics, OGB-arxiv) and heterophily graphs (Wisconsin, Cornell, Texas (Pei et al., 2020), Flickr). The results, shown in Table 5, reveal that for heterophily graphs such as Wisconsin, Cornell, and Texas, the prediction variance gap between low-degree and poisoned nodes is smaller compared to homophily graphs. However, the prediction variance of poisoned nodes remains significantly higher than that of clean nodes, even for those with low degrees. This suggests that a backdoored GNN model shows high confidence in the target class when a node is linked to a backdoor trigger, but when the trigger is removed, it shifts to high confidence in another class. In contrast, for clean nodes, edge dropping typically does not cause a significant shift in prediction confidence.
Table 5: Comparison of prediction variance between clean nodes with low degrees and poisoned nodes.
Dataset Degree-1 Degree-2 Degree-3 Target Nodes
Wisconsin 0.64 1.94 1.48 6.08
Cornell 0.79 1.33 1.27 5.16
Texas 0.65 1.48 1.28 4.10
Cora 0.51 0.66 0.56 12.70
Citeseer 0.36 0.70 0.69 5.41
Pubmed 0.67 0.80 0.57 10.67
Physics 0.44 0.52 0.36 18.80
Flickr 0.01 0.02 0.01 2.56
OGB-arxiv 0.57 0.79 0.67 10.67
Appendix G Training Algorithm
We summarize the training method of RIGBD for training a backdoor robust GNN node classifier in Algorithm 1. Specifically, we start by randomly initializing the parameters $\theta_{b}$ for an L-layer GNN node classifier $f_{b}$, which uses the graph convolution operation defined in Eq. 3 (line 1). In each iteration of the training loop (lines 2-4), we update $\theta_{b}$ by training $f_{b}$ on the backdoored graph $\mathcal{G}_{T}$ using supervised learning until the model converges. After convergence, we enter a second loop (lines 5-8) where we conduct random edge dropping on $\mathcal{G}_{T}$ and perform inference with the backdoor model $f_{b}$ (line 6). For each node, we calculate the prediction variance $s(i)$ using Eq. 2 (line 7). Nodes with prediction variance $s(i)>\tau$ are selected as candidate poisoned nodes $\mathcal{V}_{s}$, where $\tau$ is calculated by Eq. 5. Next, we randomly initialize $\theta$ for another L-layer GNN node classifier $f$ (line 10). In the final training loop (lines 11-13), we update $\theta$ by training $f$ using the robust training strategy defined in Eq. 6 until convergence. The algorithm concludes by returning the backdoor robust GNN node classifier $f$ (line 14).
@@ Appendix G Training Algorithm
1:  Randomly initialize $\theta_{b}$ for a $L$-layer GNN node classifier $f_{b}$ with graph convolution operation defined in Eq. 3;
@@
7:     Calculate the prediction variance $s(i)$ for each node by Eq. 2;
@@
9:  Select candidate poisoned nodes $\mathcal{V}_{s}$ for nodes with prediction variance $s(i)>\tau$, where $\tau$ is calculated by Eq. 5.
@@
12:     Update $\theta$ by training $f$ using the robust training strategy defined in Eq. 6;
@@
Appendix H Additional Details of Experiment Settings
H.1 Dataset statistics
Cora, Citeseer and PubMed are citation networks where nodes denote papers, and edges depict citation relationships. In Cora and Citeseer, each node is described using a binary word vector, indicating the presence or absence of a corresponding word from a predefined dictionary. In contrast, PubMed employs a TF/IDF weighted word vector for each node. For both datasets, nodes are categorized based on their respective research areas.
Coauther Physics is a co-authorship graph based on the Microsoft Academic Graph from the KDD Cup 2016 challenge (Sinha et al., 2015). Nodes are authors, that are connected by an edge if they co-authored a paper; node features represent paper keywords for each author’s papers, and class labels indicate most active fields of study for each author.
Flicker (Zeng et al., 2019): In this graph, each node symbolizes an individual image uploaded to Flickr. An edge is established between the nodes of two images if they share certain attributes, such as geographic location, gallery, or user comments. The node features are represented by a 500-dimensional bag-of-word model provided by NUS-wide. Regarding labels, we examined the 81 tags assigned to each image and manually consolidated them into 7 distinct classes, with each image falling into one of these categories.
OGB-arxiv is a citation network encompassing all Computer Science arXiv papers cataloged in the Microsoft Academic Graph. Each node is characterized by a 128-dimensional feature vector, which is derived by averaging the skip-gram word embeddings present in its title and abstract. Additionally, the nodes are categorized based on their respective research areas. The statistics details of these datasets are summarized in Table 6.
@@ H.2 Attack methods
    GTA (Xi et al., 2021): GTA utilizes a trigger generator that crafts subgraphs as triggers tailored to individual samples. The optimization of the trigger generator focuses exclusively on the backdoor attack loss, disregarding any constraints related to trigger detectability.
@@
    UGBA (Dai et al., 2023): UGBA selects representative and diverse nodes as poisoned nodes to fully utilize the attack budget. An adaptive trigger generator is optimized with a constraint loss so that the generated triggers are ensured to be similar to the target nodes.
@@
    DPGBA (Dai et al., 2023): DPGBA introduce adversarial learning strategy to generate in-distribution triggers. A novel loss is proposed to help adaptive trigger generator to generate efficient in-distribution triggers.
H.3 Compared methods
@@
    Prune (Dai et al., 2023): Prune is designed to prune edges that linking nodes with low similarity to remove those triggers that are totally different from clean nodes.
@@
    OD (Zhang et al., 2024): OD introduces graph auto-encoder and filter out nodes with high reconstruction loss to remove those triggers that are outliers compared to clean ones.
@@
    ABL (Li et al., 2021a): ABL exploits two key weaknesses of backdoor attacks: 1) models learn backdoored data much faster than clean data, with stronger attacks causing faster convergence; 2) the backdoor task is tied to a specific class. Thus, they propose a local gradient ascent (LGA) technique to trap the loss value of each example around a certain threshold. They also propose a global gradient ascent (GGA) loss function to unlearn the backdoor with a small subset of backdoor examples while continuing to learn from the remaining clean examples.
@@
    RS (Wang et al., 2021): Randomized smoothing on graphs was first proposed to counter adversarial structural perturbations. We adopt this method and set the drop ratio $\beta>0.5$ to balance the defense performance and clean accuracy. To further ensure clean accuracy, we introduce an adversarial training strategy during the training phase.
@@
    GNNGuard (Zhang & Zitnik, 2020): GNNGuard is a robust defense method for GNNs that protects against adversarial attacks by leveraging node similarity to filter out adversarial edges. It employs a multi-stages defense strategy that dynamically adjusts edge weights during training, enhancing the model’s resilience to structural perturbations.
@@
    RobustGCN (Zhu et al., 2019): RobustGCN enhances the robustness of GCN against adversarial attacks by using Gaussian distributions as node representations, which absorb the effects of adversarial changes. It also introduces a variance-based attention mechanism that assigns different weights to node neighborhoods based on their variances, reducing the propagation of adversarial effects through the graph.
H.4 Detailed comparison of RIGBD to baseline defense methods
Aside from the defense methods we mentioned, such as Prune and OD, which aim to defend against attack methods by exploiting the characteristics of generated triggers, to the best of our knowledge, our RIGBD is the first work that specifically aims to defend against dirty-label node-level graph backdoor attacks that involve generating a backdoor trigger and linking it to the target node from the perspective of training a backdoor-robust model.
Since such backdoor attack methods inherently belong to the category of poisoning attacks, we include GNNGuard and RobustGCN as baselines because they both have the ability to mitigate the influence of malicious neighbors, which may reduce the impact of backdoor triggers as these triggers are neighbors of the target nodes. However, their defense performance is not satisfactory, especially when the generated triggers are similar to the target nodes.
We also include randomized smoothing because if we can randomly drop edges with a high drop ratio during multiple inferences, we could obtain the final prediction as if the backdoor trigger were dropped. However, the clean accuracy of this method cannot be guaranteed, as too many edges are dropped during the classification task.
The core idea of ABL is that the model quickly learns backdoored data, after which a local gradient ascent technique is used to trap the loss value of each example. However, their framework requires manual tuning of the isolation ratio, and a high isolation ratio inevitably leads to a performance drop on clean data when their losses are trapped. This significantly reduces the method’s utility in real-world applications, as defenders typically do not know the number of poisoned target nodes.
In contrast, (i) Our method is theoretically guaranteed and capable of identifying those poisoned target nodes with high precision without knowledge of the number of poisoned target nodes. (ii) With high precision in detecting poisoned nodes, our framework makes only minimal changes to the graph structure of clean nodes, ensuring that the clean accuracy remains comparable to that of a model trained on a clean graph. (iii) With the proposed robust training strategy, our framework is robust to different hyperparameter selections, making it easy to implement and enhancing its utility in real-world applications.
H.5 Implementation details